Karen biked at 13 mph for a time but got a flat tire. She then walked at 5 mph. She traveled a total of 93 miles. Had she biked the whole time, she would have gone 117 miles. How many hours did she walk?
Answer by scott8148(6628)

"Had she biked the whole time, she would have gone 117 miles"
__ so the total time is 117/13 or 9hr

let x="hours walking", so 9-x=hours biking

5(x)+13(9-x)=93 __ 5x+117-13x=93 __ subtracting 117 __ -8x=-24